1.Write the equation of a line in slope-intercept form that is perpendicular to the line Y = 3/5X + -1 and passes through the point (-9, 4).

Answer by richwmiller(17219)

one problem per submission no similar problems
y=mx+b
3/5m=-1
m=-5/3
4=-5/3*(-9)+b
4=-15+b
11=b
y=-5/3x+11